OpenStudy (anonymous):

What are the exact solutions of x^2=4-7x?

OpenStudy (anonymous):

think about it yourself. All along i will be helping you. BTW, it's math. isn't it?

OpenStudy (anonymous):

ok, let's see

OpenStudy (anonymous):

x^2=4-7x tha's what you have. We want to solve for x, correct?

OpenStudy (anonymous):

x^2=4-7x to do so we move 4 - 7x on the other side and we get x^2-4+7x=0 so, what's next? what do you think?

OpenStudy (precal):

Factor if you can or use the quadratic formula